OpenStudy (anonymous):

simplify:(2x^2y^4)^3

OpenStudy (ajprincess):

(a^b)^c =a^(b*c) (2^3)*((x^2)^3)*((y^4)^3) =8x^6y^12

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

yw

OpenStudy (ajprincess):

Is that clear?

OpenStudy (anonymous):

yes